$SPAD/src/input t111293.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Coyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool t111293.output
)set message test on
)set message auto off
)clear all

———-

This file tests things fixed on or about Nov 12, 1993

unions and records in function position — * —

--S 1 of 13
y := operator y
--R
--R
--R   (1)  y
--R                                                          Type: BasicOperator
--E 1
--S 2 of 13
deq := differentiate(y x, x, 2) + differentiate(y x, x) + y x
--R
--R
--R         ,,       ,
--R   (2)  y  (x) + y (x) + y(x)
--R
--R                                                    Type: Expression(Integer)
--E 2
--S 3 of 13
solve(deq, y, x).basis
--R
--R
--R                       x     x
--R               +-+   - -   - -      +-+
--R             x\|3      2     2    x\|3
--R   (3)  [cos(-----)%e   ,%e   sin(-----)]
--R               2                    2
--R                                              Type: List(Expression(Integer))

--E 3
)clear all

———-

Other objects in function position — * —

--S 4 of 13
f := sin
--R
--R
--R   (1)  sin
--R                                                          Type: Variable(sin)
--E 4
--S 5 of 13
f 5
--R
--R
--R   (2)  sin(5)
--R                                                    Type: Expression(Integer)
--E 5
--S 6 of 13
f 5.6
--R
--R
--R   (3)  - 0.6312666378 7232131147
--R                                                                  Type: Float
--E 6
--S 7 of 13
g(f,x) == f x
--R
--R                                                                   Type: Void
--E 7
--S 8 of 13
g(cos, x)
--R
--R   Compiling function g with type (Variable(cos),Variable(x)) ->
--R      Expression(Integer)
--R
--R   (5)  cos(x)
--R                                                    Type: Expression(Integer)
--E 8
--S 9 of 13

g(f, x)
--R
--R   Compiling function g with type (Variable(sin),Variable(x)) ->
--R      Expression(Integer)
--R
--R   (6)  sin(x)
--R                                                    Type: Expression(Integer)
--E 9
--S 10 of 13
g(log, 8.38)
--R
--R   Compiling function g with type (Variable(log),Float) -> Float
--R
--R   (7)  2.1258479144 939916724
--R                                                                  Type: Float
--E 10
)clear all
--S 11 of 13
sin := [1,2,3,4,5,6,7]
--R
--R
--R   (1)  [1,2,3,4,5,6,7]
--R                                                  Type: List(PositiveInteger)
--E 11
--S 12 of 13
sin 4
--R
--R
--R   (2)  4
--R                                                        Type: PositiveInteger
--E 12
--S 13 of 13
sin(4)$Expression(Integer)
--R
--R
--R   (3)  sin(4)
--R                                                    Type: Expression(Integer)
--E 13
)spool
)lisp (bye)

———-

References^{[1] nothing}

6